Systems Staff Meeting Notes 02/16/04

Systems Staff Meeting Notes
February 16, 2004

Present: Rich Bennett, Cynthia Bowker, Will Chaney, Bill Covey, Debra Fetzer, Winston Harris, Bobbie Meng, Hilton Soto, Geoffrey Stewart, Ying Tang
Announcements
  1. Cynthia may be getting a queen-sized bed.
  2. Bobbie found a 24K gold ring (like Ying’s) at the beach this weekend.
  3. Today is the Postal Worker’s Revenge day (aka President’s Day) – no mail or banking.
  4. Yesterday’s weather was gorgeous (finally!).
Review Minutes of 2/9/04
  1. Tiered Response Priorities – will be posted on Systems Web page.
Moves/Construction Status
  1. TNT Recap – most of the group went out to BDF (ALF) on Friday to tour the facility and witness the shelving of books. It houses a huge pile of books. The lighting is horrible. It would not be a bad idea to wear a hardhat while shelving out there – the shelving is so high that if a book were to drop, you could be seriously injured. In addition to the books moving there from West, there is an ongoing reclass project – the 900’s from Dewey are either going to storage (ALF) or Library of Congress. That makes for a large number of books at ALF - and there is two year’s worth of new material still to come.
    Access Services is not accepting new receipts until March – they will take them at ALF and put them in the area where they are charging the electrical vehicle.
    Van problems – we have one library passenger van here on campus. They rest are housed out at ALF. Fleet coordination is to be coordinated by Russ in Facilities. There are no available vans with which to commute.
    Special Collections windows – they need auxiliary lighting – it’s very dark in there. The windows are supposed to be fixed by February 20 – this Friday.
    The Library now has advertising space on the moving marquee signs on campus.
    Public Services Division – grad student orientation is up by 300-500%. They recently got 3-5 people in attendance, as opposed to 0-2.
    Library East – patron usage is way down. Most are coming in only to check out videos, much to the dismay of the staff. Public workstations are running at only 50% utilization. On the other hand, MSL is crazy. HSS&S staff, however, are not participating over there.
  2. One more port in ALF was activated on Friday.
  3. West – Dee Hawes is the only staff member still working in West. She has a student workstation set up down the hallway that is utilized once a week. Bobbie had the machines moved closer. Ben Walker may still have a workstation setup – not sure. Outside of the book movers, no other staff members are working in West. Circulation is working at ALF.
Ongoing Systems Updates
Equipment Received/Installed/Repaired
  1. Dell GX110’s and GX200’s currently in the vault are going to HSS&S East room 1. The E4200’s from HSS&S are going to Education.
  2. 26 barcode readers incoming - all can read UF GatorOne cards. We’re still seeing lots of abuse with the barcode readers.
Liaisons
  1. Will sent out an email Friday listing the WSIDs that are slated for domain removal - not responding to the virus scan. He has received 12-15 replies. There are still 113 that cannot be scanned and will be deleted. Liaisons have been given one week to fix the problematic machines. The primary things they should look for are: computer browser and server service is started, file and print sharing is enabled (checked) and the Administrators group contains the domain administrators group. Some machines have simply been powered off and are sitting around. We need to pick those machines up. Geoffrey notes that 3Com Nics will power off when not active – need to disable power management. We have a lot of 3Com Nics out there. Will has taken note of that fact.
Networking
  1. Email server problem this morning – Exchange C drive has less than 10MB space. Will has been deleting old virus files – there were over 200MB on Friday. The Ref Express mailbox has been sitting untouched for 2½ years – had 28,000 virus messages – Bill’s machine did not have enough memory to delete them all at once. Old viruses are being wrapped into emails and sent out. There were two new variants to MyDoom last week. AOL and Hotmail may or may not make it to/from campus.
  2. Server room renovation – AC is tentatively set to start tomorrow. Will take a total of 3 days: installation, electric and inspection. Power is complete; network ports are in the ceiling. Facilities want to move the security servers this week, but next week works better.
  3. CNS is scheduled to terminate the extra drop in the Grand Reading Room today.
Software Projects Status
  1. Marc updates for Betsy Simpson are complete.
  2. Current Net batch uploads now being tested.
  3. Storage Retrieval – Denise Bennett is reporting some problems between the staff and patron side – may be retaining sign-ons. Winston is testing it.
  4. Map/GIS – meeting with Erich last week. Ying finished changes, now working on new ones for Search.
  5. Winston meeting at Tigert – accounts receivable transfers – looks easy to get from file to PeopleSoft. There could be a potential problem getting out of ALEPH and into PeopleSoft – would require a translation utility.
  6. ALEPH migration - Grace and Winston had a meeting with Lee about getting auxiliary data to PeopleSoft – six-character input field may need to expand. ILLIAD will stay, Hilton’s software program must move to PeopleSoft.
Web Projects
  1. New UF Web page is up.
Student and Department Happenings
  1. Nikia is back for the rest of the semester.


Back to the Systems Homepage

Last updated February 23, 2004
by Debra Fetzer
debharr@mail.uflib.ufl.edu